On January 1, 2016, Dreamworld Co. began construction of a new warehouse. The building was finished and ready for use on September 30, 2017. Expenditures on the project were as follows:

January 1, 2016 $300,000

September 1, 2016 $450,000

December 31, 2016 $450,000

March 31, 2017 $450,000

September 30, 2017 $300,000

Dreamworld had $5,000,000 in 12% bonds outstanding through both years.

Dreamworld's average accumulated expenditures for 2016 was:

a.

$300,000

b.

$450,000

c.

$525,000

d.

$600,000
Answer #1

Dreamworld's average accumulated expenditures for 2016 was:

January 1, 2016 = $300,000 x 12/12 = $300,000
September 1, 2016 = $450,000 x 4/12 = $150,000
December 31, 2016 = $450,000 x 0/12 = $0

Total = $300000 + $150000 = $450000

Option (b) is correct